Could somone help me with this? I can't figure it out.
The sum of two numbers is 376. The first number is 182 less than the second number. What are the two numbers?

Answer by scott8148(6628)

"The sum of two numbers is 376" ___ a + b = 376

"The first number is 182 less than the second number" ___ a = b - 182

substituting ___ (b - 182) + b = 376 ___ 2b - 182 = 376